Ticket Number: HEA-NOC/20020910-01 Ticket Status: UPDATE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Dublin PoP
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020911 0800 UTC+1
Ticket Closed: Problem End: 20020911 1000 UTC+1
Repatching HEAnet server and office LANs and migrating to new
Access to and from HEAnet servers will be lost for approximately 30
seconds as ethernet links negotiotiate and ARP caches timeout.
20020911 0835 cw all switchovers complete.
20020911 1030 cw problems with ip routing function, reverted at 0940
server lan still affect, working to resolve problem.
Time to Fix:
Will monitor for next 24 hours.